OpenStudy (anonymous):

can someone please HELP write a equation for each line. x-instecpt -2 y-intercept -6 HELP

OpenStudy (anonymous):

It goes through the point \((-2,0)\) and \((0,-6)\) Can't you use two points to make the line?

OpenStudy (anonymous):

The formula for two points is: \[ y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1) \]

OpenStudy (anonymous):

\[ \frac{y-y_1}{x-x_1}=\frac{y_2-y_1}{x_2-x_1}=\frac{\Delta y}{\Delta x}=m \]A fundamental property of lines is that the slope is some constant.

OpenStudy (anonymous):

\[2x+y+6=0\]